OpenStudy (anonymous):

can someone help me...i think i'm forgetting something really simple here... integrate 1/(x^2)

jimthompson5910 (jim_thompson5910):

1/(x^2) = x^(-2)

jimthompson5910 (jim_thompson5910):

from there, you just use the power rule

jimthompson5910 (jim_thompson5910):

which is int(x^n dx) = (1/(n+1)) * x^(n+1) + C

OpenStudy (anonymous):

omg...i am actually embarrassed! of course...geez...thanks jim!

jimthompson5910 (jim_thompson5910):

no worries, happens to me all the time you're welcome

OpenStudy (anonymous):

here's a medal! haha take care!

jimthompson5910 (jim_thompson5910):

oh I should mention that there's a special case if n = -1, then it won't work because you'll have n+1 = 0 in the denominator on the right side but it turns out that the integral of 1/x or x^(-1) is just ln(x)+C, which isn't so bad